OpenStudy (superhelp101):

Which of the following changes requires a reducing agent? Zn yields Zn2+ MnO2 yields Mn2+ 2Br- yields Br2 SO2 yields SO42-

OpenStudy (superhelp101):

@aaronq ?

OpenStudy (aaronq):

Reducing means GAINING electrons, as a result the oxidation state/number becomes more negative.

OpenStudy (superhelp101):

oh so if it is "gaining electrons is it Zn yields Zn2+ ?

OpenStudy (aaronq):

recall that cations eg \(Zn^{2+}\) have lost electrons.

OpenStudy (aaronq):

anions have gotten more eg. \(Cl^-\)

OpenStudy (aaronq):

In order to do this question, you need to assign oxidation states to all the atoms.

OpenStudy (superhelp101):

isn't Zn is being oxidized from 0 to +2 and thus requires a reducing agent

OpenStudy (aaronq):

nope, if the question was asking "which one requires an oxidizing agent" then that would be the answer.

OpenStudy (superhelp101):

hmm oh okay so how do i know which changes requires a reducing agent ?

OpenStudy (aaronq):

you need to determine which of these goes from a higher (more positive) oxidation state to a lower (more negative) oxidation state. So, it boils down to assigning oxidation states to all the atoms.

OpenStudy (superhelp101):

SO2 yields SO42- okay is this one right because it goes from a higher other words more positive oxidation to a low more negative oxidation ?

OpenStudy (aaronq):

whats the oxidation state of S on each side of the equation?

OpenStudy (superhelp101):

O

OpenStudy (superhelp101):

+2 +6 ?? i really don't know :/

OpenStudy (aaronq):

SO2 Oxygen is (almost) always -2, because there are 2 of them and the compound is neutral, Sulfur has to be +4 to balance out the charges. -4+x=0 x=+4 SO42- here there are 4 oxygens, each one -2, but there is also a negative charge on the compound. -8+x=-2 x=+6 Sulfur went from +4 to +6

OpenStudy (aaronq):

Try \(\sf MnO_2 \rightarrow Mn^{2+}\)

OpenStudy (superhelp101):

MnO2 --- +4 Mn2+ ---- +2

OpenStudy (aaronq):

yep! you should specify that those are the oxidation states of Manganese, but i get what you mean. Now assess, did Mn get oxidized or reduced?

OpenStudy (superhelp101):

reduced

OpenStudy (aaronq):

yep! thats it

OpenStudy (superhelp101):

Omg thank you soo much!!

OpenStudy (aaronq):

no problem! here are a couple of rules and examples when assigning oxidation states, Rules (most often true, except in rare occasions): Group 1 always +1 Group 2 always +2 Oxygen always -2 Group 7 always -1 Noble gases, elemental or homonuclear (the same element) molecules are zero The rest you have to figure out by subtracting or adding to achieve whatever charge the compound bears. eg. H2O H is group 1, so H = +1, there are 2 of them O must balance the charge so, O=-2 HCl H=+1, Cl=-1 MgSO4, you have to know that SO4^2- is a polyatomic ion, treat that as it's own separate thing: O=-2 Total contribution of O=-2(4) = -8 ^4 oxygen atoms Charge on ion=-2, so S has to fulfil the rest which is 6, so S=+6 back to MgSO4, so we treat SO4 as entity and it bears a -2 charge, Mg therefore has to be +2 to balance the charges.

OpenStudy (superhelp101):

oh ok! Thanks again! :D

OpenStudy (aaronq):

no probs!